### v1

@@ Appendix B Experimental details
Baseline algorithms. In Section 7, we compared our immax algorithm with well-known baselines, including the cross-entropy (ce) loss, Re-Weighting (rw) method (Xie and Manski, 1989; Morik et al., 1999), Balanced Softmax (bs) loss (Jiawei et al., 2020), Equalization loss (Tan et al., 2020), Logit Adjusted (la) loss (Menon et al., 2021), Class-Balanced (cb) loss (Cui et al., 2019), the focal loss in (Ross and Dollár, 2017) and the ldam loss in (Cao et al., 2019).
@@
  • Re-Weighting (rw) method (Xie and Manski, 1989; Morik et al., 1999): Each sample is re-weighted by the inverse of its class’s sample size and subsequently normalized such that the average weight within each mini-batch is 1. This is equivalent to minimizing the loss function given below:
@@
  • focal loss in (Ross and Dollár, 2017):
@@
  • equal loss: following (Tan et al., 2020), $p$ is chosen from
    $\left\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\right\}$
    and $\lambda$ is chosen from
    $\left\{0.176,0.5,0.8,1.5,1.76,2.0,3.0,5.0\right\}\times 10^{-3}.$
  • la loss: following (Menon et al., 2021), $\tau$ is chosen from
    $\left\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0\right\}$
    and
    $\left\{1.5,2.0,2.5,3.0,3.5,4.0,4.5,5.0,5.5,6.0,6.5,7.0,7.5,8.0,8.5,9.0,9.5,10.0\right\}.$
    When $\tau=1$ (the suggested value in (Menon et al., 2021)), the la loss is equivalent to the bs loss. We observed improved performance for small values of $\tau<1$ when minimizing the standard zero-one misclassification loss. Therefore, we conducted a finer search between 00 and $1$.
  • cb loss: following (Cui et al., 2019), $\gamma$ is chosen from
    $\left\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,0.99,0.999,0.9999\right\}.$
    While the default values of $\left\{0.9,0.99,0.999,0.9999\right\}$ are suggested in (Cui et al., 2019), we observed that they are not effective for minimizing the standard zero-one misclassification loss. We found that performance is typically better when $\gamma$ is close to 00.
  • focal loss: $\gamma$ is chosen from
    $\left\{1.0,1.5,2.0,2.5,3.0,3.5,4.0,4.5,5.0,5.5,6.0,6.5,7.0,7.5,8.0,8.5,9.0,9.5,10.0\right\}$
    and
    $\left\{0.0,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\right\}$
    following (Ross and Dollár, 2017). We observe that performance is typically better when $\gamma$ is less than $1$. Therefore, we conducted a finer search between 00 and $1$.
  • ldam loss: following (Cao et al., 2019), $C$ is chosen from
    $\left\{10^{-4},10^{-3},10^{-2},10^{-1},1.0,10.0,100.0,1000.0,10000.0\right\}$
    and
    $\left\{5\times 10^{-4},5\times 10^{-3},5\times 10^{-2},5\times 10^{-1},5.0,50.0,500.0,5000.0\right\}.$
  • immax loss: following Section 4 and Appendix F.4, $\rho_{k}$ is searched in the range
    $\left[\frac{m_{k}^{\frac{1}{3}}}{\sum_{j\in[c]}m_{j}^{\frac{1}{3}}}-5,\frac{m_{k}^{\frac{1}{3}}}{\sum_{j\in[c]}m_{j}^{\frac{1}{3}}}+5\right]$
    with a step size of 1. In the step imbalanced setting, we assign identical $\rho_{k}$ values to minority classes and distinct $\rho_{k}$ values to frequent classes before the search.
@@
which can be achieved by any all-positive $h^{*}_{{\mathsf{L}}_{c_{+},c_{-}}}$ such that $h^{*}_{{\mathsf{L}}_{c_{+},c_{-}}}(x)\geq 0$. Thus, $h^{*}_{{\mathsf{L}}_{c_{+},c_{-}}}$ differs from $h^{*}_{\ell_{0-1}}$, which implies that ${\mathsf{L}}_{c_{+},c_{-}}$ is not Bayes-consistent with respect to $\ell_{0-1}$.
@@ Appendix D Binary Classification: Proofs
Proof D.1.
When $yh(x)\leq 0$, we have $\Phi_{\rho_{+}}(yh(x))=\Phi_{\rho_{-}}(yh(x))=1$, so the equality holds. When $yh(x)>0$, we have $y>0\iff h(x)>0$ and $y<0\iff h(x)<0$, which also implies the equality.
@@
$\operatorname*{\mathbb{E}}_{y}\left[\ell_{0-1}(h,x,y)\mid x\right]$ $=\eta(x)\mathds{1}_{h(x)<0}+\left(1-\eta(x)\right)\mathds{1}_{h(x)\geq 0}$
@@
which completes the proof.
@@
where the last equality stems from the fact that the variables $\sigma_{i}$ and $-\sigma_{i}$ are distributed in the same way. This proves the first inequality. The second inequality, can be derived in the same way by using the second inequality of (Mohri et al., 2018, Theorem 3.3).
@@ D.4 Uniform Margin Bound for Imbalanced Binary Classification
$\operatorname*{\mathbb{P}}\bigg{[}\sup_{\begin{subarray}{c}h\in{\mathscr{H}}\\ \rho_{+}\in(0,r_{+}]\\ \rho_{-}\in(0,r_{-}]\end{subarray}}{\mathscr{R}}_{\ell_{0-1}}(h)-\widehat{\mathscr{R}}_{S}^{\rho_{+},\rho_{-}}(h)-4\mathfrak{R}_{m}^{\rho_{+},\rho_{-}}({\mathscr{H}})-\sqrt{\frac{\log\log_{2}(2r_{+}/\rho_{+})}{m}}-\sqrt{\frac{\log\log_{2}(2r_{-}/\rho_{-})}{m}}-\epsilon>0\bigg{]}$
$\qquad\leq 4e^{-2m\epsilon^{2}},$
where we used the fact that $\mathfrak{R}_{m}^{\rho_{+}/2,\rho_{-}/2}({\mathscr{H}})=2\mathfrak{R}_{m}^{\rho_{+},\rho_{-}}({\mathscr{H}})$. This proves the first statement. The second statement can be proven in a similar way.
@@ D.5 Linear Hypotheses
The first inequality makes use of the Cauchy-Schwarz inequality and the bound on $\left\|w\right\|$, the second follows by Jensen’s inequality, the third by $\operatorname*{\mathbb{E}}[\sigma_{i}\sigma_{j}]=\operatorname*{\mathbb{E}}[\sigma_{i}]\operatorname*{\mathbb{E}}[\sigma_{j}]=0$ for $i\neq j$, the fourth by $\sup_{i\in I_{+}}\left\|x_{i}\right\|=r_{+}$ and $\sup_{i\in I_{-}}\left\|x_{i}\right\|=r_{-}$, and the last one by $\left\|x_{i}\right\|\leq r$.
@@ Appendix E Extension to Multi-Class Classification
In this section, we extend the previous analysis and algorithm to multi-class classification. We will adopt the same notation and definitions as previously described, with some slight adjustments. In particular, we denote the multi-class label space by ${\mathscr{Y}}=[c]\coloneqq\left\{1,\ldots,c\right\}$ and a hypothesis set of functions mapping from ${\mathscr{X}}\times{\mathscr{Y}}$ to $\mathbb{R}$ by ${\mathscr{H}}$. For a hypothesis $h\in{\mathscr{H}}$, the label ${\sf h}(x)$ assigned to $x\in{\mathscr{X}}$ is the one with the largest score, defined as ${\sf h}(x)=\operatorname*{argmax}_{y\in{\mathscr{Y}}}h(x,y)$, using the highest index for tie-breaking. For a labeled example $(x,y)\in{\mathscr{X}}\times{\mathscr{Y}}$, the margin $\rho_{h}(x,y)$ of a hypothesis $h\in{\mathscr{H}}$ is given by $\rho_{h}(x,y)=h(x,y)-\max_{y^{\prime}\neq y}h(x,y^{\prime})$, which is the difference between the score assigned to $(x,y)$ and that of the next-highest scoring label. We define the multi-class zero-one loss function as $\ell^{\rm{multi}}_{0-1}\coloneqq\mathds{1}_{{\sf h}(x)\neq y}$. This is the target loss of interest in multi-class classification.
@@ E.1 Multi-Class Imbalanced Margin Loss
The main margin bounds in this section are expressed in terms of this loss function. The parameters $\rho_{k}>0$, for $k\in[c]$, represent the confidence margins imposed by a hypothesis $h$ for instances labeled $k$. The following result provides an equivalent expression for the class-imbalanced margin loss function. The proof is included in Appendix F.1.
Lemma E.2.
${\mathsf{L}}_{{\boldsymbol{\rho}}}(h,x,y)=\sum_{k=1}^{c}\Phi_{\rho_{k}}\left(\rho_{h}(x,y)\right)1_{{\sf h}(x)=k}.$
@@ E.2 ${\mathscr{H}}$-Consistency
Theorem E.3 (${\mathscr{H}}$-Consistency bound for multi-class class-imbalanced margin loss).
The proof is included in Appendix F.2. The next section presents generalization bounds based on the empirical multi-class class-imbalanced margin loss, along with the ${\boldsymbol{\rho}}$-class-sensitive Rademacher complexity and its empirical counterpart defined below. Given a sample $S=\left(x_{1},\ldots,x_{m}\right)$, for any $k\in[c]$, we define $I_{k}=\left\{i\in\left\{1,\ldots,m\right\}\mid y_{i}=k\right\}$ and $m_{k}=|I_{k}|$ as the number of instances labeled $k$.
@@ Appendix F Multi-Class Classification: Proofs
F.1 Proof of Lemma E.2
See E.2
Proof F.1.
When $\rho_{h}(x,y)\leq 0$, we have $\Phi_{\rho_{k}}\left(\rho_{h}(x,y)\right)=1$ for any $k\in[c]$, so the equality holds. When $\rho_{h}(x,y)>0$, we have $y=k\iff\rho_{h}(x,k)>0\iff{\sf h}(x)=k$, which also implies the equality.
F.2 Proof of Theorem E.3
See E.3
Proof F.2.
@@
$\operatorname*{\mathbb{E}}_{y}\left[\ell^{\rm{multi}}_{0-1}(h,x,y)\mid x\right]$ $=\sum_{y\in{\mathscr{Y}}}p(y\!\mid\!x)\mathds{1}_{{\sf h}(x)\neq y}=1-p({\sf h}(x)\!\mid\!x),$
@@
$\operatorname*{\mathbb{E}}_{y}\left[\ell^{\rm{multi}}_{0-1}(h,x,y)\mid x\right]-\inf_{h\in{\mathscr{H}}}\operatorname*{\mathbb{E}}_{y}\left[\ell^{\rm{multi}}_{0-1}(h,x,y)\mid x\right]=\max_{y\in{\mathscr{Y}}}p(y\!\mid\!x)-p({\sf h}(x)\!\mid\!x).$
@@
$=1-p({\sf h}(x)\!\mid\!x)+p({\sf h}(x)\!\mid\!x)\max\left(0,1-\frac{\rho_{h}(x,{\sf h}(x))}{\rho_{{\sf h}(x)}}\right)$
$=1-p({\sf h}(x)\!\mid\!x)\min\left(1,\frac{\rho_{h}(x,{\sf h}(x))}{\rho_{{\sf h}(x)}}\right).$
@@
$\operatorname*{\mathbb{E}}_{y}\left[{\mathsf{L}}_{{\boldsymbol{\rho}}}(h,x,y)\mid x\right]-\inf_{h\in{\mathscr{H}}}\operatorname*{\mathbb{E}}_{y}\left[{\mathsf{L}}_{{\boldsymbol{\rho}}}(h,x,y)\mid x\right]$ $=\max_{y\in{\mathscr{Y}}}p(y\!\mid\!x)-p({\sf h}(x)\!\mid\!x)\min\left(1,\frac{\rho_{h}(x,{\sf h}(x))}{\rho_{{\sf h}(x)}}\right)$
$\geq\max_{y\in{\mathscr{Y}}}p(y\!\mid\!x)-p({\sf h}(x)\!\mid\!x)$
@@
which completes the proof.
@@
This proves the second inequality. The first inequality, can be derived in the same way by using the first inequality of (Mohri et al., 2018, Theorem 3.3).

### v2

@@ Appendix B Experimental details
Baseline algorithms. In Section 7, we compared our immax algorithm with well-known baselines, including the cross-entropy (ce) loss, Re-Weighting (rw) method (Xie & Manski, 1989; Morik et al., 1999), Balanced Softmax (bs) loss (Jiawei et al., 2020), Equalization loss (Tan et al., 2020), Logit Adjusted (la) loss (Menon et al., 2021), Class-Balanced (cb) loss (Cui et al., 2019), the focal loss in (Ross & Dollár, 2017) and the ldam loss in (Cao et al., 2019).
@@
  • Re-Weighting (rw) method (Xie & Manski, 1989; Morik et al., 1999): Each sample is re-weighted by the inverse of its class’s sample size and subsequently normalized such that the average weight within each mini-batch is 1. This is equivalent to minimizing the loss function given below:
@@
  • focal loss in (Ross & Dollár, 2017):
@@
  • equal loss: following (Tan et al., 2020), $p$ is chosen from $\left\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\right\}$ and $\lambda$ is chosen from $\left\{0.176,0.5,0.8,1.5,1.76,2.0,3.0,5.0\right\}\times 10^{-3}$.
  • la loss: following (Menon et al., 2021), $\tau$ is chosen from $\left\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0\right\}$ and $\left\{1.5,2.0,2.5,3.0,3.5,4.0,4.5,5.0,5.5,6.0,6.5,7.0,7.5,8.0,8.5,9.0,9.5,10.0\right\}$. When $\tau=1$ (the suggested value in (Menon et al., 2021)), the la loss is equivalent to the bs loss. We observed improved performance for small values of $\tau<1$ when minimizing the standard zero-one misclassification loss. Therefore, we conducted a finer search between 00 and $1$.
  • cb loss: following (Cui et al., 2019), $\gamma$ is chosen from $\left\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,0.99,0.999,0.9999\right\}$. While the default values of $\left\{0.9,0.99,0.999,0.9999\right\}$ are suggested in (Cui et al., 2019), we observed that they are not effective for minimizing the standard zero-one misclassification loss. We found that performance is typically better when $\gamma$ is close to 00.
  • focal loss: $\gamma$ is chosen from $\left\{1.0,1.5,2.0,2.5,3.0,3.5,4.0,4.5,5.0,5.5,6.0,6.5,7.0,7.5,8.0,8.5,9.0,9.5,10.0\right\}$ and $\left\{0.0,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\right\}$ following (Ross & Dollár, 2017). We observe that performance is typically better when $\gamma$ is less than $1$. Therefore, we conducted a finer search between 00 and $1$.
  • ldam loss: following (Cao et al., 2019), $C$ is chosen from $\left\{10^{-4},10^{-3},10^{-2},10^{-1},1.0,10.0,100.0,1000.0,10000.0\right\}$ and $\left\{5\times 10^{-4},5\times 10^{-3},5\times 10^{-2},5\times 10^{-1},5.0,50.0,500.0,5000.0\right\}$.
  • immax loss: following Section 4 and Appendix F.4, each $\rho_{k}$ is searched within a symmetric interval centered at the theoretically optimal value $\rho_{k}^{*}=\frac{m_{k}^{\frac{1}{3}}}{\sum_{j\in[c]}m_{j}^{\frac{1}{3}}}$, with width $\delta_{k}$; that is, over the interval $\left[\rho_{k}^{*}-\delta_{k},\rho_{k}^{*}+\delta_{k}\right]$, where $\delta_{k}$ is chosen relative to the scale of $\rho_{k}^{*}$. In particular, we set $\delta_{k}\approx 0.8\rho_{k}^{*}$ and evaluate approximately 10 values within each interval. Empirically, performance is not sensitive to variations within this neighborhood. In the step imbalanced setting, we assign identical $\rho_{k}$ values to minority classes and distinct $\rho_{k}$ values to frequent classes before the search.
@@
which can be achieved by any all-positive $h^{*}_{{\mathsf{L}}_{c_{+},c_{-}}}$ such that $h^{*}_{{\mathsf{L}}_{c_{+},c_{-}}}(x)\geq 0$. Thus, $h^{*}_{{\mathsf{L}}_{c_{+},c_{-}}}$ differs from $h^{*}_{\ell_{0-1}}$, which implies that ${\mathsf{L}}_{c_{+},c_{-}}$ is not Bayes-consistent with respect to $\ell_{0-1}$. ∎
@@ Appendix D Binary Classification: Proofs
Proof.
When $yh(x)\leq 0$, we have $\Phi_{\rho_{+}}(yh(x))=\Phi_{\rho_{-}}(yh(x))=1$, so the equality holds. When $yh(x)>0$, we have $y>0\iff h(x)>0$ and $y<0\iff h(x)<0$, which also implies the equality. ∎
@@
$\operatorname*{\mathbb{E}}_{y}\left[\ell_{0-1}(h,x,y)\mid x\right]$ $=\eta(x)1_{h(x)<0}+\left(1-\eta(x)\right)1_{h(x)\geq 0}$
@@
which completes the proof. ∎
@@
where the last equality stems from the fact that the variables $\sigma_{i}$ and $-\sigma_{i}$ are distributed in the same way. This proves the first inequality. The second inequality, can be derived in the same way by using the second inequality of (Mohri et al., 2018, Theorem 3.3). ∎
@@ D.4 Uniform Margin Bound for Imbalanced Binary Classification
$\operatorname*{\mathbb{P}}\bigg{[}$ $\sup_{\begin{subarray}{c}h\in{\mathscr{H}}\\ \rho_{+}\in(0,r_{+}]\\ \rho_{-}\in(0,r_{-}]\end{subarray}}{\mathscr{R}}_{\ell_{0-1}}(h)-\widehat{\mathscr{R}}_{S}^{\rho_{+},\rho_{-}}(h)-4\mathfrak{R}_{m}^{\rho_{+},\rho_{-}}({\mathscr{H}})-\sqrt{\frac{\log\log_{2}(2r_{+}/\rho_{+})}{m}}-\sqrt{\frac{\log\log_{2}(2r_{-}/\rho_{-})}{m}}-\epsilon>0\bigg{]}\leq 4e^{-2m\epsilon^{2}},$
where we used the fact that $\mathfrak{R}_{m}^{\rho_{+}/2,\rho_{-}/2}({\mathscr{H}})=2\mathfrak{R}_{m}^{\rho_{+},\rho_{-}}({\mathscr{H}})$. This proves the first statement. The second statement can be proven in a similar way. ∎
@@ D.5 Linear Hypotheses
The first inequality makes use of the Cauchy-Schwarz inequality and the bound on $\left\|w\right\|$, the second follows by Jensen’s inequality, the third by $\operatorname*{\mathbb{E}}[\sigma_{i}\sigma_{j}]=\operatorname*{\mathbb{E}}[\sigma_{i}]\operatorname*{\mathbb{E}}[\sigma_{j}]=0$ for $i\neq j$, the fourth by $\sup_{i\in I_{+}}\left\|x_{i}\right\|=r_{+}$ and $\sup_{i\in I_{-}}\left\|x_{i}\right\|=r_{-}$, and the last one by $\left\|x_{i}\right\|\leq r$. ∎
@@ Appendix E Extension to Multi-Class Classification
In this section, we extend the previous analysis and algorithm to multi-class classification. We will adopt the same notation and definitions as previously described, with some slight adjustments. In particular, we denote the multi-class label space by ${\mathscr{Y}}=[c]\coloneqq\left\{1,\ldots,c\right\}$ and a hypothesis set of functions mapping from ${\mathscr{X}}\times{\mathscr{Y}}$ to $\mathbb{R}$ by ${\mathscr{H}}$. For a hypothesis $h\in{\mathscr{H}}$, the label ${\mathsf{h}}(x)$ assigned to $x\in{\mathscr{X}}$ is the one with the largest score, defined as ${\mathsf{h}}(x)=\operatorname*{argmax}_{y\in{\mathscr{Y}}}h(x,y)$, using the highest index for tie-breaking. For a labeled example $(x,y)\in{\mathscr{X}}\times{\mathscr{Y}}$, the margin $\rho_{h}(x,y)$ of a hypothesis $h\in{\mathscr{H}}$ is given by $\rho_{h}(x,y)=h(x,y)-\max_{y^{\prime}\neq y}h(x,y^{\prime})$, which is the difference between the score assigned to $(x,y)$ and that of the next-highest scoring label. We define the multi-class zero-one loss function as $\ell^{\rm{multi}}_{0-1}\coloneqq 1_{{\mathsf{h}}(x)\neq y}$. This is the target loss of interest in multi-class classification.
@@ E.1 Multi-Class Imbalanced Margin Loss
The main margin bounds in this section are expressed in terms of this loss function. The parameters $\rho_{k}>0$, for $k\in[c]$, represent the confidence margins imposed by a hypothesis $h$ for instances labeled $k$. The following result provides an equivalent expression for the class-imbalanced margin loss function. The proof is included in Appendix F.1. See 5.2
@@ E.2 ${\mathscr{H}}$-Consistency
See 5.3 The proof is included in Appendix F.2. The next section presents generalization bounds based on the empirical multi-class class-imbalanced margin loss, along with the ${\boldsymbol{\rho}}$-class-sensitive Rademacher complexity and its empirical counterpart defined below. Given a sample $S=\left(x_{1},\ldots,x_{m}\right)$, for any $k\in[c]$, we define $I_{k}=\left\{i\in\left\{1,\ldots,m\right\}\mid y_{i}=k\right\}$ and $m_{k}=|I_{k}|$ as the number of instances labeled $k$.
@@ Appendix F Multi-Class Classification: Proofs
F.1 Proof of Lemma 5.2
See 5.2
Proof.
When $\rho_{h}(x,y)\leq 0$, we have $\Phi_{\rho_{k}}\left(\rho_{h}(x,y)\right)=1$ for any $k\in[c]$, so the equality holds. When $\rho_{h}(x,y)>0$, we have $y=k\iff\rho_{h}(x,k)>0\iff{\mathsf{h}}(x)=k$, which also implies the equality. ∎
F.2 Proof of Theorem 5.3
See 5.3
Proof.
@@
$\operatorname*{\mathbb{E}}_{y}\left[\ell^{\rm{multi}}_{0-1}(h,x,y)\mid x\right]$ $=\sum_{y\in{\mathscr{Y}}}p(y\!\mid\!x)1_{{\mathsf{h}}(x)\neq y}=1-p({\mathsf{h}}(x)\!\mid\!x),$
@@
$\operatorname*{\mathbb{E}}_{y}\left[\ell^{\rm{multi}}_{0-1}(h,x,y)\mid x\right]-\inf_{h\in{\mathscr{H}}}\operatorname*{\mathbb{E}}_{y}\left[\ell^{\rm{multi}}_{0-1}(h,x,y)\mid x\right]=\max_{y\in{\mathscr{Y}}}p(y\!\mid\!x)-p({\mathsf{h}}(x)\!\mid\!x).$
@@
$=1-p({\mathsf{h}}(x)\!\mid\!x)+p({\mathsf{h}}(x)\!\mid\!x)\max\left(0,1-\frac{\rho_{h}(x,{\mathsf{h}}(x))}{\rho_{{\mathsf{h}}(x)}}\right)$
$=1-p({\mathsf{h}}(x)\!\mid\!x)\min\left(1,\frac{\rho_{h}(x,{\mathsf{h}}(x))}{\rho_{{\mathsf{h}}(x)}}\right).$
@@
$\operatorname*{\mathbb{E}}_{y}\left[{\mathsf{L}}_{{\boldsymbol{\rho}}}(h,x,y)\mid x\right]-\inf_{h\in{\mathscr{H}}}\operatorname*{\mathbb{E}}_{y}\left[{\mathsf{L}}_{{\boldsymbol{\rho}}}(h,x,y)\mid x\right]$ $=\max_{y\in{\mathscr{Y}}}p(y\!\mid\!x)-p({\mathsf{h}}(x)\!\mid\!x)\min\left(1,\frac{\rho_{h}(x,{\mathsf{h}}(x))}{\rho_{{\mathsf{h}}(x)}}\right)$
$\geq\max_{y\in{\mathscr{Y}}}p(y\!\mid\!x)-p({\mathsf{h}}(x)\!\mid\!x)$
@@
which completes the proof. ∎
@@
This proves the second inequality. The first inequality, can be derived in the same way by using the first inequality of (Mohri et al., 2018, Theorem 3.3). ∎